Graph the line that passes though the given point and has the given slope. (4,2), slope - 1/5
Answer by mananth(16946)

(4,2), slope - 1/5
y=mx+b is the equation of a line
equation ---
2 = (-1/5)*4+c
2=-4/5 +c
c=2+4/5
=14/5
y=(-1/5)x+14/5